OpenStudy (anonymous):

4x^2-x-3=0 solve by completing the square x= ??

OpenStudy (anonymous):

x=-0.75

OpenStudy (anonymous):

4x2 - x - 3 = 0 ⇒ x=1 , x=-0.75

OpenStudy (zarkon):

\[4x^2-x-3=4\left(x-\frac{1}{8}\right)^{2}-\frac{49}{16}\]

jimthompson5910 (jim_thompson5910):

4x^2-x-3=0 4(x^2-x/4-3/4)=0 4(x^2-x/4+1/16-1/16-3/4)=0 4((x-1/8)^2-1/64-3/4)=0 4((x-1/8)^2-49/64)=0 4(x-1/8)^2-4(49/64)=0 4(x-1/8)^2-49/16=0 4(x-1/8)^2=49/16 (x-1/8)^2=49/64 x-1/8 = sqrt(49/64) or x-1/8=-sqrt(49/64) x-1/8 = 7/8 or x-1/8 = -7/8 x = 7/8+1/8 or x = -7/8+1/8 x = 8/8 or x = -6/8 x = 1 or x = -3/4 Answers are \[\large x=1 \textrm{ or } x=-\frac{3}{4}\]

OpenStudy (anonymous):

jim, you are good!!!